# Mental Disorder Classification via Temporal Representation of Text

### v1

@@ Abstract
Mental disorders pose a global challenge, aggravated by the shortage of qualified mental health professionals. Mental disorder prediction from social media posts by current LLMs is challenging due to the complexities of sequential text data and the limited context length of language models. Current language model-based approaches split a single data instance into multiple chunks to compensate for limited context size. The predictive model is then applied to each chunk individually, and the most voted output is selected as the final prediction. This results in the loss of inter-post dependencies and important time variant information, leading to poor performance. We propose a novel framework which first compresses the large sequence of chronologically ordered social media posts into a series of numbers. We then use this time variant representation for mental disorder classification. We demonstrate the generalization capabilities of our framework by outperforming current SOTA in three different mental conditions: depression, self-harm, and anorexia by an absolute improvement of 5% in F1 score. We investigate the situation where current data instances fall within the context length of language models and present empirical results highlighting the importance of temporal properties of textual data. Furthermore, we utilize the proposed framework for a cross-domain study, exploring commonalities across disorders and the possibility of inter-domain data usage.
@@ 1 Introduction
To handle these issues, we must account for two aspects: 1) preserving the post-identity and order and 2) global classification i.e. classifying the entire sequence of posts together instead of classifying chunks of social media text. Using current LLMs in this scenario can be challenging due to computational costs and the high context length of social media posts. For instance, the average context length of a subject is 11.6K words (max 89.4K) Table 1, which exceeds the input context length of current LLMs.
Hence, the chunking-based methods can not address these concerns, while other LLM-based approaches may prove infeasible within the given context. This motivates the need for a representation technique capable of compressing information while retaining the time-variant properties of social media posts. We build upon these insights and propose a framework to account for these issues. We perform rigorous analyses to establish the importance of temporal properties in mental disorders (Section 7.3). We also discuss the performance of the current language models when the input length is within the capacity of the language model (Section 8). We open-source our code for academic purposes [^2] GitHub link is redacted. Our contributions are:
  • A first-of-its-kind representation method which transforms textual data from social media posts into a time series format to capture the time-dependent patterns of a patient (Figure 2). This provides a compressed representation of the textual data while reducing the floating point operation by at least 330 times (Table 3) compared to SOTA.
  • A novel technique incorporating temporal data for mental disorder identification by using foundational deep learning models. Our approach outperforms language model-based methods by 5% in F1 score (Table 2) across three mental conditions: anorexia, depression, and self-harm.
  • A cross-domain study (Section 9) of the three disorders to understand the commonality across disorders. We investigate the possibility of cross-domain data usage (Table 9), which can further benefit the identification of low-resource mental disorders.
@@ 2 Related Work
A few studies closely align with our approach to constructing temporal representations of social media posts. Reece et al. (2017) employed state-space temporal analysis using day and week as the time window for depression detection. De Choudhury et al. (2013) examined a user’s tweets within a single day to derive various behavioural measures and constructed a time series for each measure. Chen et al. (2020) created a time series representation of the mood profile using traditional sentiment retrieval models. A significant limitation of these approaches is their reliance on low-level features, which do not provide a deeper semantic understanding of the emotional aspects of human language.
@@
In this work, we propose a novel post-representation method which captures the temporal patterns of an individual while compressing a large amount of textual information.
@@ 4.1 Anchor Embedding Generation
The initial step involves creating an anchor embedding which serves as a semantic anchor of a specific disorder’s condition class.
@@ 4.3 Time Series Classification
For time series classification (Figure 2 part (3)), we explore two approaches: (i) Feature extraction-based, where we train a feed-forward network on statistical and temporal features extracted from the time series data, using the top 30 features for classification. (ii) Representation learning-based, where we directly feed the raw time series data as input to the model. For representation learning-based approaches, we use three different architectures, namely: 1) LSTM Yu et al. (2019) 2) 1D CNN Tang et al. (2020), and 3) Time-Series Transformer Zerveas et al. (2021). We use the method of threshold moving Zhang et al. (2020) to decide the best probability threshold for the decision boundary to account for data imbalance (Table 1). We describe all the experimental details including features list in the Appendix C.
@@ 5 Baseline Approaches
Additionally, we use MPNet Song et al. (2020) as a baseline in both zero-shot and fine-tuned settings. To compare the results with current SOTA language models, we report results on GPT3.5 turbo Brown et al. (2020) and MentalLLaMA-chat-13B Yang et al. (2023b) instruction tuned on large mental health data. We follow Aragon et al. (2023) for evaluation and training. All the baseline methods used the chunking (k=35) and majority voting-based approach. The F1 score of the positive class serves as our standard metric for assessing the experimental results. We provide a detailed explanation for baselines in Appendix B and present experimental results on additional baselines in Table 6, Appendix C.1.
@@ 6 Results and Observations
The poor performance of GPT-3.5-turbo and MentalLLaMA-chat-13B (Table 2) may be due to the chunking strategy, which does not consider the longitudinal history of the users’s texts and removes the inter-post dependencies. It is also unlikely that a single chunk has sufficient information for accurate mental disorder prediction. We provide additional insights with an example in Appendix F. To go one step further in this analysis, we present the results on the instances when the text is within the range of maximum input size in Section 8.
@@ 7 Analysis
Our approach demonstrates significant improvement in self-harm and depression prediction while showing a competitive performance in anorexia prediction (Table 2). This highlights the necessity of considering the complete context of the input data and its temporal information. Our methodology differs from other approaches that focus on understanding the absolute semantics of a subject. Instead, we model the semantic contrast between the posts of the subject and the anchor embedding representing the condition class. This approach enhances our ability to identify disorders characterized by high semantic contrast, like self-harm.
@@ 7.2 Efficiency Analysis
Table 3: Total number of Floating points operation required for a single forward pass. Here, "Anor", "Depr", and "SH" stand for anorexia, depression, and self-harm. Models marked with * are baselines.
To study the computational efficiency of our framework, we report the number of floating point operations (Flos) in a single forward pass Kaplan et al. (2020). We observe that the maximum number of Flos in the proposed methodology is 25.5 million, while the minimum number of Flos in the considered baselines is 8.42 billion. This reduces the total number of floating point operations by 330 times in the worst-case scenario (refer Appendix E).
Figure 4: Results for temporal analysis: F1 scores comparison between the permuted input data and the ordered input data for three disorders
@@ 7.3 Temporal Analysis
To understand the contributions of temporal properties of the data towards the performance, we perform the training after permuting the input data five times in random order. This ensures that the post-wise temporal property of data is lost. From Figure 4, we observe a significant dip in performance as compared to our original setup and note that the performance is worse or comparable to Aragon et al. (2023). These results empirically establish the importance of post-wise temporal properties of social media data.
@@ 8 Full Context Analysis
In this section, we compare our approach with LLMs in the scenario when the data instances have context length within the capacity of the LLM. The main purpose of this experiment is to understand if LLMs can understand and capture the temporality directly from the natural language. We perform this experiment using MentalLLaMA Yang et al. (2023b) because of its superior performance in mental disorder identification over GPT and other LLMs. For this experiment, we only consider the data instances with word count within the capacity of MentalLLaMA (i.e. 2048 tokens). We report the data statistics in the Table 8. It can be noted that we have a very small number of instances of depression; therefore, the results may not be conclusive for this case. We report all the results in Figure 5. We observe that our approach outperforms MentalLLaMA in all three tasks. This observation aligns with recent studies like Liu et al. (2024), which demonstrates the inability of LLMs to utilize long context inputs.
Figure 5: F1 scores over the condition class in three eRisk tasks by considering up to 2k context length.
9 Cross-domain Study
@@
To understand the commonality of the disorder $\mathcal{D}_{1}$ with respect to disorder $\mathcal{D}_{2}$, we take the cosine similarity between the anchor embedding of the $\mathcal{D}_{1}$ disorder with the post embeddings of $\mathcal{D}_{2}$ disorder arranged in chronological order. After obtaining this series of cosine similarities (Section 4), we use it as the train set and evaluate the test set of the disorder $\mathcal{D}_{1}$. This step aims to extract shared information about the mental disorder under consideration from the data of other disorders. For example, to study the linguistic similarity of depression with self-harm, we take the cosine similarity between the anchor embedding of depression and social media posts embedding from self-harm. Using the acquired series as the training set, we report the test set results from depression. We observe that the results are significantly better than the random baseline and competitive with the SOTA in certain cases. We present results on all three pairs (6 combinations) in Table 9, Appendix H. Specifically, we observe that anorexia and self-harm show good cross-domain F1 scores, whereas the other pairs involving depression show sub-optimal results as compared to SOTA. This may be due to the extra-linguistic features in the depressed class subjects, which may act as noise for the other disorders. While a substantial number of patients with anorexia or self-harm also experience depression, not every individual with depression exhibits symptoms of these mental disorders. Overall, this indicates that linguistic cues essential for classifying one disorder may be present in others, hinting at the potential of leveraging data for other domains.
@@ 10 Conclusion and Future Work
This work proposes a novel framework to incorporate temporal representation of textual data for the identification of Anorexia, Depression, and Self-harm from social media data. Our methodology utilizes fundamental deep-learning architecture and surpasses LLM-based baselines by accounting for temporality and the full context of the input data. Our cross-domain analysis highlights the overlapping linguistic cues among the disorders and hints at the possibility of leveraging data from different mental disorders.
Our work can be extended by exploring more complex mental disorders such as schizophrenia, personality disorders, bipolar disorder etc. The task of mental disorder classification can be enhanced by including other modalities like audio and visual signals that give insights into behavioural patterns. The cross-domain study of disorders also encourages the exploration of multitask setups for mental disorder prediction.
@@
Our study’s limitations arise from the challenges inherent in analyzing mental disorders through social media data. Notably, as discussed in our error analysis, the presence of out-of-context posts and individuals’ reluctance to openly express their mental health challenges due to stigma can lead to false negatives. Additionally, the incomplete context within social media posts, where some content is removed or censored for community well-being, can hinder the framework’s accuracy in identifying mental disorders, emphasizing the complexities of using online data.
@@
When conducting an analysis of social media content pertaining to mental disorders, it is essential to address valid concerns regarding individual privacy and ethical considerations. These concerns stem from handling sensitive and personal information, including discussions about emotions and health-related issues. All the examples of social media posts mentioned in this paper are hypothetical representations of the actual data as we do not have permission to publish any portion of the dataset (e.g. example post) other than summary statistics. It is worth noting that we have exclusively utilized publicly available datasets, specifically Reddit datasets (Section 3.1) and eRisk collections (Section 4.1). We have diligently adhered to the terms of use and user agreements associated with these collections. Furthermore, the datasets we have employed are anonymized, and our research does not involve any direct interaction with social media users. Given these conditions and practices, our study does not necessitate review and approval by an Ethics Committee Board.

### v2

@@ Abstract
Mental disorders pose a global challenge, aggravated by the shortage of qualified mental health professionals. Mental disorder prediction from social media posts by current LLMs is challenging due to the complexities of sequential text data and the limited context length of language models. Current language model-based approaches split a single data instance into multiple chunks to compensate for limited context size. The predictive model is then applied to each chunk individually, and the most voted output is selected as the final prediction. This results in the loss of inter-post dependencies and important time variant information, leading to poor performance. We propose a novel framework which first compresses the large sequence of chronologically ordered social media posts into a series of numbers. We then use this time variant representation for mental disorder classification. We demonstrate the generalization capabilities of our framework by outperforming current SOTA in three different mental conditions: depression, self-harm, and anorexia by an absolute improvement of 5% in F1 score. We also investigate the situation when current data instances fall within the context length of language models and present empirical results highlighting the importance of temporal properties of textual data. Furthermore, we utilize the proposed framework for a transfer-learning study, exploring commonalities across disorders and the possibility of inter-domain data usage.
@@ 1 Introduction
To handle these issues, we must account for two aspects: 1) preserving the post-identity and order and 2) global classification i.e. classifying the entire sequence of posts together instead of classifying chunks of social media text. Using current LLMs here can be challenging due to computational costs and the high context length of social media posts. For instance, the average context length of a subject is 11.6K words (max 89.4K) Table 1, which exceeds the input context length of current LLMs.
Hence, the chunking-based methods can not address these concerns, while other LLM-based approaches may prove infeasible within the given context. This motivates the need for a representation technique capable of compressing information while retaining the time-variant properties of social media posts. We build upon these insights and propose a framework to account for these issues. We perform rigorous analyses to establish the importance of temporal properties in mental disorders (Section 7.3). We also discuss the performance of the current language models when the input length is within the capacity of the language model (Section 8). Our source code is available on GitHub [^2] GitHub Link for academic purposes.
Our contributions are:
  • A representation method for mental-health domain which compresses textual data from social media posts into a time series format to capture the time-dependent patterns of a patient (Figure 2). This provides a temporal representation of the textual data while reducing the floating point operations (FLOPs) by at least 330 times (Table 3) compared to SOTA.
  • A novel framework incorporating temporal data for mental disorder identification by using foundational deep learning models. Our approach outperforms language model-based methods by 5% in F1 score (Table 2) across three mental conditions: anorexia, depression, and self-harm.
  • A transfer-learning study (Section 9) of the three disorders to understand the commonality across disorders. We investigate the possibility of cross-domain data usage (Table 5), which can further benefit the identification of low-resource mental disorders.
In the subsequent sections, we discuss recent works that are relevant to our contributions. We then discuss the dataset used and the proposed method, followed by the insights from our experiments.
@@ 2 Related Work
A few studies closely align with our approach to constructing temporal representations of social media posts. Reece et al. (2017) employed state-space temporal analysis using day and week as the time window for depression detection. De Choudhury et al. (2013) examined a user’s tweets within a single day to derive various behavioural measures and constructed a time series for each measure. Chen et al. (2020) created a time series representation of the mood profile using traditional sentiment retrieval models. Lee et al. (2023) proposed a multi-task framework aimed at predicting suicidality for bipolar disorder patients by categorizing individual posts into different suicidality levels. Sawhney et al. (2021b) introduced graph-based approaches to model user interaction and the temporality of posts, while Sawhney et al. (2021a) modeled "phases" at the post level, requiring post-level annotations for suicide ideation detection. A significant limitation of these approaches is either their reliance on low-level features, which fail to capture the deeper semantic understanding of emotional aspects in human language, or the need for user interaction data and post-level annotations, which are not available in our setup.
@@
In this work, we propose a novel framework which captures the temporal patterns of social media posts without the use of majority voting and chunking.
@@ 4.1 Anchor Embedding Generation
The initial step involves creating an anchor embedding which serves as a semantic anchor of a specific disorder’s condition class. A similar approach was used by Fei and Liu (2015) to address covariate shifts in social media text classification.
@@ 4.3 Time Series Classification
For time series classification (Figure 2 part (3)), we explore two approaches: (i) Feature extraction-based, where we train a feed-forward network on statistical and temporal features extracted from the time series data, using the top 30 features for classification. This selection of raw time series features was motivated by our interest in assessing the performance of a basic feature engineering approach for this task. We extracted these features to capture relevant characteristics from the data. (ii) Representation learning-based, where we directly feed the raw time series data as input to the model. For representation learning-based approaches, we use three different architectures, namely: 1) LSTM Yu et al. (2019) 2) 1D CNN Tang et al. (2020), and 3) Time-Series Transformer Zerveas et al. (2021). We use the method of threshold moving Zhang et al. (2020) to decide the best probability threshold for the decision boundary to account for data imbalance (Table 1). We describe all the experimental details including features list in the Appendix C.
@@ 5 Baseline Approaches
Additionally, we use MPNet Song et al. (2020) as a baseline in both zero-shot and fine-tuned settings. To compare the results with current SOTA language models, we report results on GPT3.5 turbo Brown et al. (2020) and MentalLLaMA-chat-13B Yang et al. (2024) instruction tuned on large mental health data. We follow Aragon et al. (2023) for evaluation and training. All the baseline methods used the chunking (k=35) and majority voting-based approach. The F1 score of the positive class serves as our standard metric for assessing the experimental results. We provide a detailed explanation for baselines in Appendix B and present experimental results on additional baselines in Table 9, Appendix C.1.
@@ 6 Results and Observations
The poor performance of GPT-3.5-turbo and MentalLLaMA-chat-13B (Table 2) may be due to the chunking strategy, which does not consider the longitudinal history of the users’s texts and removes the inter-post dependencies. Since important signals are concentrated in some of the posts, most of the posts may not have sufficient information for accurate mental disorder prediction. We provide additional insights with an example in Appendix F. To go one step further in this analysis, we present the results on the instances when the text is within the range of maximum input size in Section 8.
@@ 7 Analysis
Our approach demonstrates improvement in self-harm and depression prediction while showing a competitive performance in anorexia prediction (Table 2). This highlights the necessity of considering the complete context of the input data and its temporal information. Our methodology differs from other approaches that focus on understanding the absolute semantics of a subject. Instead, we model the semantic contrast between the posts of the subject and the anchor embedding representing the condition class. This approach enhances our ability to identify disorders characterized by high semantic contrast, like self-harm.
@@ 7.2 Efficiency Analysis
Table 3: Total number of FLOPs required for a single forward pass. Here, "Anor", "Depr", and "SH" stand for anorexia, depression, and self-harm. Models marked with * are baselines.
To study the computational efficiency of our framework, we report the number of floating point operations (FLOPs) in a single forward pass Kaplan et al. (2020). We observe that the maximum number of FLOPs in the proposed methodology is 25.5 million, while the minimum number of FLOPs in the considered baselines is 8.42 billion. This reduces the total number of floating point operations by 330 times in the worst-case (Appendix E).
Figure 4: Results for temporal analysis: F1 scores comparison between the permuted input data and the ordered input data for three disorders of condition class
@@ 7.3 Temporal Analysis
Temporal analysis is essential in our setup since anchor embedding was derived from condition class users. The cosine similarity values for condition classes are expected to be higher than the control class in the time-series representation (Figure 3). To understand if our model is picking up the temporal order (not just the magnitude-related information), we perform the training after permuting the input data five times in random order. This ensures that the post-wise temporal property of data is lost. From Figure 4, we observe a significant dip in performance as compared to our original setup and note that the performance is worse or comparable to Aragon et al. (2023). These results empirically establish the importance of post-wise temporal properties of social media data.
7.4 Ablation Study
In this section, we study the importance of anchor embedding by introducing three different ablation setups without an anchor:
Direct Encoding: This involves direct encoding of the posts by using MPNet Agarwal et al. (2024) and utilizes the complete 768 dimension vector without anchor embedding.
Probabilities Values: This involves calculating probability values using three Huggingface models, each fine-tuned for specific mental health conditions: a BERT Devlin et al. (2019) model for self-harm [^3] Self-Harm-HF, and RoBERTa Liu et al. (2019) models for anorexia [^4] Anorexia-HF and depression [^5] Depression-HF. The probabilities are then arranged in chronological order for time series classification, without creating anchors.
Plutchik-wheel-based Emotions: This setup involves using eight different emotions (anger, anticipation, joy, trust, fear, surprise, sadness and disgust) based on Plutchik-wheel Sawhney et al. (2021a) for individual posts. Here, we used the probability scores of a post for each emotion and performed multivariate time series classification in a similar setup.
Model F1 P R
Anorexia
Direct Encoding 0.68 0.63 0.80
Probabilities Values 0.76 0.79 0.73
Plutchik-wheel-based Emotions 0.30 0.20 0.72
LSTM (Anchor-based Approach) 0.79 0.84 0.74
Depression
Direct Encoding 0.66 0.64 0.70
Probabilities Values 0.47 0.46 0.48
Plutchik-wheel-based emotions 0.62 0.69 0.56
LSTM (Anchor-based Approach) 0.75 0.79 0.71
Self-Harm
Direct Encoding 0.62 0.72 0.56
Probabilities Values 0.73 0.81 0.66
Plutchik-wheel-based emotions 0.47 0.37 0.67
LSTM (Anchor-based Approach) 0.83 0.93 0.75
Table 4: Ablation Study - Comparison of three alternative setups without anchor embedding
We observe that the scores for various ablated model variants are consistently lower than those of our proposed anchor-based method, underscoring the importance of anchor embeddings.
@@ 8 Full Context Analysis
In this section, we compare our approach with LLMs in the scenario when the data instances have context length within the capacity of the LLM. The main purpose of this experiment is to understand if LLMs can understand and capture the temporality directly from the natural language. We perform this experiment using MentalLLaMA Yang et al. (2024) because of its superior performance in mental disorder identification over GPT and other LLMs. For this experiment, we only consider the data instances with word count within the capacity of MentalLLaMA (i.e. 2048 tokens). We report the data statistics in the Table 10. It is important to note that users in the dataset whose histories fit within the context window may also systematically differ from the larger population, which could affect the generalizability of the results. Additionally, we have a very small number of instances of depression, so the results may not be conclusive for this case. We report all the results in Figure 5. We observe that our approach outperforms MentalLLaMA in all three tasks. This observation aligns with recent studies like Liu et al. (2024), which demonstrates the inability of LLMs to utilize long context inputs.
Figure 5: F1 scores of the condition class in three eRisk tasks by considering up to 2k context length.
9 Transfer-learning Setting
@@
Model Train+Val Test F1 P R
Anorexia (A)
DisorBERT A A 0.83 0.82 0.85
LSTM A A 0.79 0.84 0.74
LSTM D A 0.75 0.68 0.83
LSTM SH A 0.80 0.85 0.75
Depression (D)
DisorBERT D D 0.69 0.56 0.89
LSTM D D 0.75 0.79 0.71
LSTM A D 0.63 0.86 0.50
LSTM SH D 0.63 0.73 0.56
Self-Harm (SH)
DisorBERT SH SH 0.74 0.72 0.76
LSTM SH SH 0.83 0.93 0.75
LSTM A SH 0.78 0.85 0.72
LSTM D SH 0.69 0.65 0.77
Table 5: Results for transfer-learning evaluations for all six combinations of disorders. Here, ‘A’ is anorexia, ‘D’ is depression and ‘SH’ is self-harm.
To understand the commonality of the disorder $\mathcal{D}_{1}$ with respect to disorder $\mathcal{D}_{2}$, we take the cosine similarity between the anchor embedding of the $\mathcal{D}_{1}$ disorder with the post embeddings of $\mathcal{D}_{2}$ disorder arranged in chronological order. After obtaining this series of cosine similarities (Section 4), we use it as the train set and evaluate the test set of the disorder $\mathcal{D}_{1}$. This step aims to extract shared information about the mental disorder under consideration from the data of other disorders. For example, to study the linguistic similarity of depression with self-harm, we take the cosine similarity between the anchor embedding of depression and social media posts embedding from self-harm. Using the acquired series as the training set, we report the test set results from depression. We observe that the results are significantly better than the random baseline and competitive with the SOTA in certain cases. We present results on all three pairs (6 combinations) in Table 5. Specifically, we observe that anorexia and self-harm show good F1 scores in the transfer-learning setting, whereas the other pairs involving depression show sub-optimal results as compared to SOTA. This may be due to the extra-linguistic features in the depressed class subjects, which may act as noise for the other disorders. While a substantial number of patients with anorexia or self-harm also experience depression, not every individual with depression exhibits symptoms of these mental disorders. Overall, this indicates that linguistic cues essential for classifying one disorder may be present in others, hinting at the potential of leveraging data for other domains.
@@ 10 Conclusion and Future Work
This work proposes a novel framework to incorporate temporal representation of textual data for the identification of Anorexia, Depression, and Self-harm from social media data. Our methodology utilizes fundamental deep-learning architecture and surpasses LLM-based baselines by accounting for temporality and the full context of the input data. Our transfer-learning analysis highlights the overlapping linguistic cues among the disorders and hints at the possibility of leveraging data from different mental disorders.
Our work can be extended by exploring more complex mental disorders such as schizophrenia, personality disorders, bipolar disorder etc. The task of mental disorder classification can be enhanced by including other modalities like audio and visual signals that give insights into behavioural patterns. The proposed framework can be also extended beyond the mental health domain in scenarios which require a temporal understanding of natural language data.
@@
Our study’s limitations arise from the challenges inherent in analyzing mental disorders through social media data. Notably, as discussed in our error analysis, the presence of out-of-context posts and individuals’ reluctance to openly express their mental health challenges due to stigma can lead to false negatives. Additionally, the incomplete context within social media posts, where some content is removed or censored for community well-being, can hinder the framework’s accuracy in identifying mental disorders, emphasizing the complexities of using online data. Other limitations of this work may arise from the use of traditional deep learning architectures. For instance, the feed-forward and CNN networks were sensitive to the choice of hyperparameters, especially for anorexia. To handle this, an extensive grid search was performed.
@@
When conducting an analysis of social media content pertaining to mental disorders, it is essential to address valid concerns regarding individual privacy and ethical considerations. These concerns stem from handling sensitive and personal information, including discussions about emotions and health-related issues. All the examples of social media posts mentioned in this paper are paraphrased and anonymized representations of the actual data as we do not have permission to publish any portion of the dataset (e.g. example post) other than summary statistics. It is worth noting that we have exclusively utilized publicly available datasets, specifically Reddit datasets (Section 3.1) and eRisk collections (Section 4.1). We have diligently adhered to the terms of use and user agreements associated with these collections. Furthermore, the datasets we have employed are anonymized, and our research does not involve any direct interaction with social media users. Given these conditions and practices, our study does not necessitate review and approval by an Ethics Committee Board.